OpenStudy (anonymous):

Write the plane on parametric form: 2x-4y+z=0 What is the general algorithm for solving a "problem" of this sort?

OpenStudy (anonymous):

Argh! This is a huge blockage of going further in my course! :(

OpenStudy (apoorvk):

Hmm. Just like any other thing. So, 3 variables, hence you need two parameters let x=p, and let y=q so, z= 4y-2z hence, (x,y,z)=(p,q,4y-2z) Get how it's done?

OpenStudy (anonymous):

Not really :( I am aware that I need two parameters. Is there an unique solution for the equation on parametric form?

OpenStudy (anonymous):

If so, the answer is supposed to be: x=-s y=t z=2s+4t

OpenStudy (anonymous):

I got no clue how they arrived at that though.

OpenStudy (anonymous):

a plane can be defined like the following: \[r = a + t_1b + t_2 c\] we have 2x-4y+z=0 i want parameters that will give me a point where this is the case so if we rearrange: z = 4y - 2x now choose parameters carefully so that this is the case: let \[y = t_2 \text{ , } x = t_1\] and using the earlier equation \[z = 4t_2 - 2t_1\] we now need \[r = a + t_1 b + t_2 c = \left(\begin{matrix}t_1 \\ t_2 \\4 t_2 - 2t_1\end{matrix}\right)\] can you see that this is a possible solution? \[a= \left(\begin{matrix}0 \\ 0 \\ 0\end{matrix}\right)\] \[b = \left(\begin{matrix}1 \\ 0 \\-2\end{matrix}\right)\] \[c = \left(\begin{matrix}1 \\ 0 \\ 4\end{matrix}\right)\]

OpenStudy (anonymous):

how was that?

OpenStudy (anonymous):

Extremely dedicated reply and easy to understand - thanks once again @eigenschmeigen !

OpenStudy (anonymous):

;D its all in a days work for your local neighbourhood eigenschmeigen!